# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ 4 Nondeterministic Polynomial-time Problem Challenge
Figure 5: Modules in NPPC
We introduce the Nondeterministic Polynomial Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. There are three main components in NPPC (as displayed in Figure 5): i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances and verify the solution with any levels of complexities, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate the users to evaluate their own models and iii) npeval, which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over problems, the number of tokens, the “aha moments”, the reasoning and solution errors, providing the in-depth analysis of the LLMs’ reasoning capabilities.
@@ 4.1 Problem Suite: npgym
Figure 6: Interaction loop between the LLM and the nygym.
Interaction Protocol.
[…]
Core Problems and Extension.
There are 25 typical NPC problems implemented in npgym. Among all NPC problems, we select a representative subset of 12 NPC problems as the core problems, ranging from the most famous NPC problems, e.g., 3SAT, Vertex Cover, and Clique, to the mathematical programming and string processing. The other 13 problems are categorized as the extension problems. A full list of the 25 problems is displayed in Table 2.
@@
Generation and Verification.
Specifically, for each problem, npgym implements two functions:
@@
Difficulty Levels.
Different NPC problems exhibit distinct combinatorial structures and computational characteristics. To establish a standardized metric for quantifying the computational complexity of these problems, npgym implements the difficulty levels [7, 13]. Each difficulty level corresponds to a specific parameterization. The current implementation of npgym stratifies each NPC problem into approximately 10 discrete difficulty levels, calibrated to induce a monotonically decreasing performance curve in LLMs, ranging from >90% success rate at level 1 to <10% at level 10. For each difficulty level, infinite number of problem instances can be generated, thus npgym is unhackable. The difficulty levels are empirically calibrated against current LLM performances and higher difficulty levels can be seamlessly included when the LLMs evolve to be more capable, which ensures that npgym is uncrushable. The full list of the difficulty levels is displayed in Appendix C.1.
@@ 4.2 Solver Suite: npsolver
Prompt Template.
The prompt template for LLMs is designed to be simple without any problem-specific knowledge and consistent across all problems. Therefore, the prompt template includes: i) problem description, which provides the concise definition of the NPC problem, including the problem name, the input and the question to be solved, ii) the context examples, where each example is formed by the instance and its corresponding solution, demonstrating the input and output patterns to help LLMs to generate the solution, iii) the target instance to solve, and iv) the general instruction about the solution format, where the solution is required to be in the JSON format for easy extracting and analyzing. We note that the structural output in JSON format may bring difficulties for LLMs to generate the correct solution, especially for the offline models, which will be analyzed in the experiments. The complete prompt template is displayed in Appendix D.
Completion with LLMs.
To streamline response extraction across various LLMs, we present npsolver, a solver suite that provides a unified interface for both online (API-based) and offline (locally deployed) models. npsolver includes: i) prompt generation, which constructs problem-specific prompts dynamically using the designed prompt templates, ii) LLM completion, that handles response generation via either online APIs supported through LiteLLM [5], or offline models via vLLM [19]; iii) solution extraction, which applies regular expressions to parse JSON-formatted responses, ensuring a consistent validation pipeline across all models; iv) error reporting, that standardizes error messages. Through the unified interface, npsolver enables both online and offline models to share a common workflow for completion and further evaluation.
@@ 4.3 Evaluation Suite: npeval
A comprehensive evaluation of the LLMs over all problems and all difficulty levels is time-consuming and expensive, mainly due to several sources of the randomness: i) the randomness of the generated instances, and ii) the randomness of the LLMs’ responses [^1] Randomizing the responses, i.e., non-zero temperature, is recommended for better performance [14].. Existing benchmarks let the LLMs solve the problem instances in a dataset, e.g., 200 instances in X-Large dataset in  [20], spanning across 5 difficulty levels and each level with 40 instances. However, we want to evaluate the LLMs’ performance on each difficulty level. Inspired by rliable [2], npeval considers the aggregated performance over different independent seeds, e.g., 3, for each difficulty level, and 30 instances are generated and evaluated for each seed, where 30 samples are generally considered as the minimum number for statistical analysis [16]. This sampling strategy enables statistically sound performance aggregation while controlling for instance-specific variance under the limited budget.
Figure 7: npeval.
Inspired by rliable [2], npeval provides the four measures to evaluate the performance: inter-quantile mean (IQM), mean, median, and the optimality gap and leverages stratified bootstrap confidence intervals (SBCIs) [11, 12] with stratified sampling for aggregate performance to report the interval estimation of the performance, which is a method that can be applied to small sample sizes and is better justified than reporting sample standard deviations. npeval provides the analysis of both prompt and completion tokens of LLMs across the problems and difficulty levels. As observed in [14], there are “aha moments” in the reasoning content. Therefore, npeval also provides the analysis of the number of “aha moments” during the reasoning. npeval provides the analysis of the errors including the errors of the generated solutions, i.e., the errors returned by npgym, and the reasoning errors, i.e., the errors in the internal reasoning process of LLMs for generating the solutions. This enables identification of the failures of LLMs. The main evaluations in npgym are displayed in Figure 7.
@@ 5 Results
Figure 8: Performance over difficulty levels
Figure 9: Ranks of models over problems
@@
5.1 Analysis of Performance
The performance of online LLMs over difficulty levels is displayed in Figure 8, where all online models exhibit a decline in accuracy as difficulty levels increase across all 12 NPC problems. Take 3SAT as an example, all online models except for DeepSeek-R1 drop from $\geq 80\%$ accuracy to close to 0% at the last level, and DeepSeek-R1 shows the slowest decline but still falls to $\leq 15\%$ accuracy. All models collapse to around or even below $10\%$ accuracy at extreme difficulty confirms that NPPC is uncrushable against the SoTA LLMs and can discriminate their capabilities. One exception is Claude-3.7-Sonnet on Superstring problem, where the accuracy is still above 50% even for the level 10, while other models are all decreased into less than 20%, which demonstrates the superiority of Claude-3.7-Sonnet to deal with long contexts, where the prompts at level 10 is more than 50K [^2] We do not continually increase the difficulty of this problem as all other models are worse than 10%.. All models perform similarly on the Bandwidth problem, which may be mainly due to the fact that none of the models are familiar with this specific problem. Both o3-mini and DeepSeek-V3-2503 demonstrate superior performance to their predecessor models, o1-mini and DeepSeek-V3, respectively, validating continually improvements in both non-reasoning and reasoning LLMs.
@@ 5.1 Analysis of Performance
5.2 Analysis of Tokens and Aha Moments
Tokens.
Figure 11 displays the token utilization across models on 3SAT. Offline models (QwQ-32B, DeepSeek-R1-32B) rapidly approach maximum token limits and incorrect solutions (red) usually take more tokens that correct solutions (blue). Among online models, DeepSeek-R1 demonstrates highest consumption (10,000-20,000 tokens) for successful solutions, while o-series models exhibit significant variance with outliers exceeding 40,000 tokens at higher complexity levels. DeepSeek-R1 and o3-mini show steeper token scaling compared to o1-mini and Claude-3.7-Sonnet, indicating advanced reasoning models leverage increased token allocation for complex problem-solving. GPT-4o variants maintain relatively efficient token utilization (<2,000) across all complexities. This quantifies the computational efficiency-performance tradeoff between specialized reasoning architectures and general-purpose models. Due to the limited space, full results over all problems are in Appendix I.
Figure 11: The number of tokens of different models on 3SAT
Aha Moments.
We also provide the analysis of the aha moments (instances of insight during reasoning, marked by phrases like “wait”) in the reasoning contents of DeepSeek-R1333The reasoning contents of o1/o3-mini are not available for analysis.. The results are displayed in Figure 12. While some problems (e.g., Superstring, TSP, and Bin Packing) show that the aha moments first increase and then decrease with difficulty, others exhibit steadier increases (QDE, 3DM) or relatively stable patterns (3SAT). This diversity suggests that advanced models’ approach to increasingly complex problems depends significantly on the specific problem domain and the necessities of investigating of different NPC problems.
Figure 12: Number of aha moments in DeepSeek-R1
@@ 5.3 Analysis of Solution Errors
The solution errors of 3SAT is displayed in Figure  13. The results show that the distribution of these errors varies across models and difficulty levels. As the difficulty increases, the frequency of certain error types tends to increase as well. For QwQ-32B, JSON ERROR dominate across all levels, which is mainly due to the reasoning process is not finished when the context reaches the limits. For other models (such as GPT-4o, Claude-3.7-Sonnet, and DeepSeek-R1), problem-specific errors (3SAT ERROR 1 and 2) become more prevalent at higher difficulty levels. DeepSeek and GPT models show a more balanced distribution of error types. The o-series models demonstrate relatively better performance at intermediate difficulties but still encounter increasing problem-specific errors at the highest complexity levels. This error analysis provides valuable insights into not just whether models fail, but specifically how they fail when tackling increasingly complex NPC problems, which could inform future improvements to reasoning capabilities in LLMs.
@@ 5.4 Analysis of Reasoning Failure Cases
For Deepseek-R1, the reasoning content of the failure cases shows several reasons that lead to wrong answers. i) cascading assumptions and placeholder/guess: DeepSeek-R1 begins with a high-level approach but quickly resorts to making assumptions to derive answers without logical deduction and considering all the conditions, and finally returns a placeholder or an educated guess; ii) manual computation errors and complexity avoidance: DeepSeek-R1 uses inefficient manual calculations (prone to errors) instead of programming, skips complex steps even the reasoning is correct, and resorts to guesses to avoid effort; iii) reliance on prompt examples: DeepSeek-R1 relies heavily on the example solution, making it waste time and get distracted by verifying and editing the solution instead of solving the problem directly; iv) reasoning repetition: DeepSeek-R1 gets stuck repeating the same logic without making further progress, wasting time and tokens. We list some typical examples of failure cases of DeepSeek-R1 in Table 3, and more examples are shown in Table 20 in Appendix L. For Claude-3.7-Sonnet, its failure cases typically exhibit more concise reasoning. Claude often outlines a high-level step-by-step approach but omits detailed calculations and rigorous verification, and it relies on approximate calculations to derive a final answer, incorrectly asserting that the result has been validated. Example is shown in Table 21 in Appendix L.
6 Conclusions

### v2

@@ 4 Nondeterministic Polynomial-time Problem Challenge
Figure 4: Overview of NPPC. (a) NPPC represents the intersection of foundation models and foundation problems. (b) The three main components of NPPC: npgym (problem generation), npsolver (solution generation), and npeval (evaluation). (c) Workflow diagram: npgym configuring generators and verifiers, npsolver using LLMs to generate solutions, and npeval measuring performance metrics.
We introduce Nondeterministic Polynomial Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. There are three main components in NPPC (as displayed in Figure 4(b)): i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances and verify the solution with any levels of complexities, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate the users to evaluate their own models and iii) npeval, which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over problems, the number of tokens, the “aha moments”, the reasoning and solution errors, providing the in-depth analysis of the LLMs’ reasoning capabilities.
@@ 4.1 Problem Suite: npgym
Interaction Protocol. Typically, NPC problems are the decision problems where given the instance $I$, the answer is “Yes” or “No”. However, the LLMs may take a random guess without reasoning for the true solution (Fan et al., 2024). Therefore, we consider a more challenging setting: given the instance $I$, the LLM needs to generate the solution $s$ for the instance. This setting will enforce the LLMs to reason for the correct solutions and the NPPC needs to provide the certificate $\sigma$ to verify the solutions generated by the LLMs. npgym provides a unified interface of NPC problems to interact with LLMs. The interaction between npgym and the LLM is displayed in Figure 4(c). npgym generates the instance $I$ with the given configuration, and the LLM receives the instance and generate the solution $s$, then the solution is verified by npgym with the output $\{\textbf{true},\textbf{false}\}$. The representation of problem instances is designed to be concise and complementary to include all necessary information for the LLMs to reason for the solution.
Core Problems and Extension. There are 25 typical NPC problems implemented in npgym. Among all NPC problems, 12 typically NPC problems are selected as the core problems, chosen for their fundamental importance and broad real-world applications across domains such as logistics and routing (TSP, Hamiltonian Cycle), network optimization (Vertex Cover, Graph 3-Colourability), resource allocation (Bin Packing, 3-Dimensional Matching), automated reasoning (3SAT), computational biology (Shortest Common Superstring), and mathematical optimization (Quadratic Diophantine Equations, Minimum Sum of Squares). The other 13 problems are categorized as the extension problems, covering specialized applications in social networks, facility location, cryptography, and data mining. A full list of the 25 problems is displayed in Table 2.
@@
Generation and Verification. Specifically, for each problem, npgym implements two functions:
@@
Difficulty Levels. NPC problems exhibit distinct combinatorial structures and computational characteristics. npgym implements the difficulty levels (Cobbe et al., 2020; Fan et al., 2024) establish a standardized metric for quantifying the computational complexity. Specifically, the difficult levels are determined with a two-stage approach: First, the parameters for NPC problems are manually configured based on problem-specific insights (e.g., graph size, constraint density) by human experts. Second, to facilitate intuitive visualization and meaningful benchmarking comparisons, we empirically calibrate the difficulty scale using LLM performance data. This calibration serves purely as a post-hoc validation step to ensure the difficulty progression is interpretable for benchmark users, which does not influence problem generation or introduce circular reasoning into the evaluation methodology. The calibration confirms that higher difficulty levels (as determined by theoretical complexity) correspond to lower LLM success rates, providing an intuitive scale where level 1 problems achieve >90% success and level 10 problems achieve <10% success.The comprehensive justification of this approach is in Appendix A.6. Appendix C.1 includes full specifications of difficulty levels.
Is There a Unified Principle for Difficulty Levels of NPC Problems? Establishing a unified principle for determining difficulty levels across all NPC problems is fundamentally challenging due to inherent differences from both theoretical and practical perspectives. From the problem perspective, the structural heterogeneity of NPC problems prevents the establishment of a universal difficulty metric. While all NPC problems are polynomially reducible to each other in theory, they exhibit vastly different characteristics in practice. These differences include: i) representation complexity, i.e., problems vary in how constraints and variables are encoded (graph structures vs. logical formulas vs. numerical constraints), ii) Search space topology, i.e., some problems have smooth difficulty landscapes while others contain sharp complexity transitions. This heterogeneity means that uniform metrics—such as simple parameter counts or constraint numbers—fail to capture the true computational difficulty that emerges during actual problem-solving. From the LLM perspective, LLMs demonstrate highly variable performance across different NPC problems, and the problem instances generated should not be too easy or too difficult, which may fail to differentiate the capabilities of LLMs. Additionally, there exists no established theoretical framework for determining the upper bounds of problem difficulty that LLMs can effectively handle, making difficulty calibration necessarily empirical and problem-specific, which justifies our two-stage approach to determine the difficulty levels.
Ever-scalingness of npgym. npgym fulfills the four desiderata of ever-scalingness. Specifically, npgym can generate enormous problem instances with arbitrary difficulty levels, enabling scaling over complexity and instance to continuously differentiate the LLMs while avoiding hacking, e.g., memorization. The scalability is fundamentally grounded in the mathematical properties of NPC problems: for any fixed instance size $n$, the solution space grows exponentially. Solution verification in npgym is computationally efficient, guaranteed by the inherent properties of NP problems where candidate solutions can be verified in polynomial time. npgym supports extensible coverage through a simple interface requiring only two core functions and difficulty specifications for adding new NP(C) problems and no specialized tool is required, enabling the benchmark to expand across diverse computational domains while maintaining consistency in evaluation methodology.
@@ 4.2 Solver Suite: npsolver
nppc_template = """
# <problem_name> Problem Description:
<problem_description>
# Examples:
<in_context_examples>
# Problem to Solve:
Problem: <problem_to_solve>
# Instruction:
“‘json
{"solution": "___" }
example_and_solution = """Problem: <example_problem>
{"solution": <example_solution>}
Figure 5: Prompt Template of NPPC
Prompt Template. The prompt template for LLMs is designed to be simple without any problem-specific knowledge and consistent across all problems. Therefore, the prompt template (displayed above) includes: i) problem description, which provides the concise definition of the NPC problem, including the problem name, the input and the question to be solved, ii) the context examples, where each example is formed by the instance and its corresponding solution, demonstrating the input and output patterns to help LLMs to generate the solution, iii) the target instance to solve, and iv) the general instruction about the solution format, where the solution is required to be in the JSON format for easy extracting and analyzing. We note that the structural output in JSON format may bring difficulties for LLMs to generate the correct solution, especially for offline models (analyzed in the experiments). More details are displayed in Appendix D.
Completion with LLMs. To streamline response extraction across various LLMs, we present npsolver, a solver suite that provides a unified interface for both online (API-based) and offline (locally deployed) models. npsolver includes: i) prompt generation, which constructs problem-specific prompts dynamically using the designed prompt templates, ii) LLM completion, that handles response generation via either online APIs supported through LiteLLM (BerriAI, 2023), or offline models via vLLM (Kwon et al., 2023); iii) solution extraction, which applies regular expressions to parse JSON-formatted responses, ensuring a consistent validation pipeline across all models; iv) error reporting, that standardizes error messages. Through the unified interface, npsolver enables both online and offline models to share a common workflow for completion.
@@ 4.3 Evaluation Suite: npeval
Comprehensive LLM evaluation across all problems and difficulty levels is computationally expensive due to the randomness in instance generation and LLM responses [^1] Randomizing responses, i.e., non-zero temperature, is used for better performance (Guo et al., 2025).. While existing benchmarks evaluate LLMs on fixed datasets (e.g., 200 instances across 5 difficulty levels in  (Lin et al., 2025)), difficulty-specific performance assessment is required, thus leading to the development of npeval (as displayed in Figure 4(c)). Inspired by rliable (Agarwal et al., 2021), npeval aggregates performance across multiple independent seeds (typically 3) for each difficulty level, generating 30 instances per seed—the minimum sample size for statistical analysis. This sampling strategy enables statistically sound performance aggregation while controlling instance-specific variance within budget constraints. npeval provides four performance measures following rliable, i.e., inter-quantile mean (IQM), mean, median, and optimality gap, which employ stratified bootstrap confidence intervals (SBCIs) with stratified sampling for aggregate performance estimation, a method suitable for small sample sizes and more robust than standard deviations. We note that IQM trims extreme values and computes the interquartile mean across runs and tasks to smooth out the randomness in responses, which highlights the consistency of the performance and complements metrics like mean/median to avoid outlier skew. The framework analyzes both prompt and completion tokens across problems and difficulty levels, as well as the number “aha moments” in reasoning processes in (Guo et al., 2025). Additionally, it categorizes errors into solution errors (detected by npgym’s verification) and reasoning errors (flaws in the LLM’s internal problem-solving process). More details can be found in Appendix C.3.
@@ 5 Results
We conduct comprehensive experiments to evaluate the reasoning capabilities of state-of-the-art LLMs across the NPPC benchmark. Our evaluation encompasses 10 representative models, including two offline medium-sized reasoning models (QwQ-32B and DeepSeek-R1-32B), four online advanced non-reasoning models (GPT-4o-mini, GPT-4o, Claude-3.7-Sonnet, and two versions of DeepSeek-V3), and four online reasoning-specialized models (DeepSeek-R1, o1-mini, and o3-mini). For each problem, we generate instances across 10 difficulty levels, with each level designed to progressively challenge model capabilities through increased problem complexity. Following our rigorous sampling strategy, we evaluate each model on 90 instances per difficulty level (30 instances across 3 independent seeds) to ensure statistical reliability.
5.1 Analysis of Performance
Figure 6: Performance over difficulty levels measured by IQM
The performance of considered LLMs over difficulty levels is displayed in Figure 6, where all models exhibit a decline in accuracy as difficulty levels increase across all 12 NPC problems. Take 3SAT as an example, all online models except for DeepSeek-R1 drop from $\geq 80\%$ accuracy to close to 0% at the last level, and DeepSeek-R1 shows the slowest decline but still falls to $\leq 15\%$ accuracy. All models collapse to around or even below $10\%$ accuracy at extreme difficulty confirms that NPPC is not crushed against the SoTA LLMs and can discriminate their capabilities. One exception is Claude-3.7-Sonnet on Superstring problem, where the accuracy is still above 50% even for the level 10, while other models are all decreased into less than 20%, which demonstrates the superiority of Claude-3.7-Sonnet to deal with long contexts, where the prompts at level 10 is more than 50K [^2] We do not continually increase the difficulty of this problem as all other models are worse than 10%.. All models perform similarly on the Bandwidth problem, which may be mainly due to the fact that none of the models are familiar with this specific problem. Both o3-mini and DeepSeek-V3-2503 demonstrate superior performance to their predecessor models, o1-mini and DeepSeek-V3, respectively, validating continually improvements in both non-reasoning and reasoning LLMs.
Figure 7: Ranks of models over problems, where the x-axis represents the rank, ranging from 1 to 10, as we evaluate 10 models, and the y-axis shows the distribution of different LLMs across the ranks.
@@ 5.1 Analysis of Performance
Takeaways NPPC can successfully decrease the performances of advanced LLMs to $10\%$ DeepSeek-R1, o3-mini and Claude are the strongest LLMs across all considered NPC problems The ranks of different LLMs depend on the specific NPC problems
5.2 Analysis of Used Tokens
Figure 9 displays the token utilization across models on 3SAT. Offline models (QwQ-32B, DeepSeek-R1-32B) rapidly approach maximum token limits and incorrect solutions (red) usually take more tokens than correct solutions (blue). Among online models, DeepSeek-R1 demonstrates highest consumption (10,000-20,000 tokens) for successful solutions, while o-series models exhibit significant variance, with outliers exceeding 40,000 tokens at higher complexity levels. DeepSeek-R1 and o3-mini show steeper token scaling compared to o1-mini and Claude-3.7-Sonnet, indicating advanced reasoning models leverage increased token allocation for complex problem-solving. GPT-4o variants maintain relatively efficient token utilization (<2,000) across all complexities. This quantifies the computational efficiency-performance tradeoff between specialized reasoning architectures and general-purpose models. Similar phenomenon are also observed in the analysis of the aha moments (instances of insight during reasoning, marked by phrases like “wait”) in the reasoning contents of DeepSeek-R1333The reasoning contents of o1/o3-mini are not available for analysis.. Due to the limited space, full results of tokens over all problems and the analysis of aha moment are displayed in Appendices I and J, respectively.
Figure 9: The number of tokens of different models on 3SAT. The correct and incorrect solutions are represented as blue and red points, respectively, and the line are the average values over all instances.
Takeaways Reasoning models can solve more difficult problems by scaling up the number of tokens used The number of tokens used first increase then decrease, indicating the failure of LLM reasoning
@@ 5.3 Analysis of Solution Errors
The solution errors of 3SAT is displayed in Figure 10. The results show that the distribution of these errors varies across models and difficulty levels. As the difficulty increases, the frequency of certain error types tends to increase as well. For QwQ-32B, JSON ERROR dominate across all levels, which is mainly due to the reasoning process is not finished when the context reaches the limits (7500 Tokens). For other models (such as GPT-4o, Claude-3.7-Sonnet, and DeepSeek-R1), problem-specific errors (3SAT ERROR 1 and 2) become more prevalent at higher difficulty levels. Interestingly, we observe that some models (e.g., DeepSeek-R1-32B, DeepSeek-V3, GPT-4o and o1-mini) exhibit increasing JSON error at higher difficulty levels, similar to how humans might lose track of formatting instructions during lengthy reasoning processes on challenging problems. DeepSeek and GPT models show a more balanced distribution of error types. The o-series models demonstrate relatively better performance at intermediate difficulties but still encounter increasing problem-specific errors at the highest complexity levels. This error analysis provides valuable insights into not just whether models fail, but specifically how they fail when tackling increasingly complex NPC problems, which could inform future improvements to reasoning capabilities in LLMs.
Takeaways Offline models, e.g., QwQ-32B, and DeepSeek-V3 series models suffer the format errors The problem-specific solution errors provide the feedback for further improving the LLMs
@@ 5.4 Analysis of Reasoning Failure Cases
For Deepseek-R1, the reasoning content of the failure cases shows several reasons that lead to wrong answers. i) cascading assumptions and placeholder/guess: DeepSeek-R1 begins with a high-level approach but quickly resorts to making assumptions to derive answers without logical deduction and considering all the conditions, and finally returns a placeholder or an educated guess; ii) manual computation errors and complexity avoidance: DeepSeek-R1 uses inefficient manual calculations (prone to errors) instead of programming, skips complex steps even the reasoning is correct, and resorts to guesses to avoid effort; iii) reliance on prompt examples: DeepSeek-R1 relies heavily on the example solution, making it waste time and get distracted by verifying and editing the solution instead of solving the problem directly; iv) reasoning repetition: DeepSeek-R1 gets stuck repeating the same logic without making further progress, wasting time and tokens. We list some typical examples of failure cases of DeepSeek-R1 in Table 3, and more examples are shown in Table 22 in Appendix L. Failure cases of Claude-3.7-Sonnet typically exhibit more concise reasoning, as it often outlines a high-level step-by-step approach but omits detailed calculations and rigorous verification, and it relies on approximate calculations to derive a final answer, incorrectly asserting that the result has been validated. More examples are shown in Tables 23 and 24 in Appendix L.
5.5 Cost of Evaluations
According to the cost analysis in Table 4, we observe significant cost variations across different models when evaluated on the same benchmark tasks. With approximately 31 million prompt tokens processed, the total cost ranges from \$10.31 for GPT-4o-mini to \$522.48 for o3-mini, representing more than a 50-fold difference. Notably, reasoning-enhanced models (o1-mini, o3-mini, and DeepSeek-R1) exhibit substantially higher completion token consumption due to the generation of extensive intermediate reasoning tokens. For instance, o3-mini generates 110 million completion tokens, 11.7 times more than GPT-4o-mini, directly contributing to its elevated operational cost. In contrast, GPT-4o-mini demonstrates the best cost-effectiveness through its efficient token generation strategy, while Claude-3.7-Sonnet, despite moderate completion tokens (11.2 million), incurs a total cost of \$269.19 due to its higher pricing (\$15/MTok for completion). The DeepSeek series models show competitive pricing under the RMB pricing structure, with DeepSeek-V3 requiring only 192.41 RMB (approximately \$27). It is worth noting that compared with static benchmarks, e.g., ZebraLogic (Lin et al., 2025), evaluating on NPPC is inherently more costly due to its ever-scalingness and the rigorous evaluation protocol. These cost metrics provide crucial economic considerations for researchers in large-scale evaluations where model inference and pricing strategies directly impact project feasibility.
Table 4: Cost for online models
Model Prompt Completion Cost
GPT-4o-mini 30964144 (\$0.15/MTok) 9442548 (\$0.6/MTok) \$10.31
GPT-4o 30963606 (\$2.5/MTok) 7786156 (\$10/MTok) \$155.27
Claude-3.7-Sonnet 33799101 (\$3/MTok) 11186272 (\$15/MTok) \$269.19
DeepSeek-V3 31490957 (2RMB/MTok) 16178388 (8RMB/MTok) 192.41RMB
o1-mini 31360984 (\$1.1/MTok) 35161551 (\$4.4/MTok) \$189.21
o3-mini 31199884 (\$1.1/MTok) 110944621 (\$4.4/MTok) \$522.48
6 Limitations and Future Work
Multimodal NP Problems. The first limitation of this work is only text-based NPC problems are considered. Extending NPPC to the multimodal domains represents a promising direction. Games like StarCraft II, Minesweeper, Pokemon and Super Mario Bros (Aloupis et al., 2015), could form the foundation of a multimodal version of NPPC. However, extending NPC problems to the multimodal domain presents significant challenges that require careful consideration and novel approaches. Two primary obstacles emerge in this endeavor: first, not all NPC problems are inherently suitable for multimodal representation, as demonstrated by problems like 3SAT which are fundamentally symbolic and lack natural visual components; second, maintaining the scalable difficulty characteristics essential to NPC problems becomes complex when incorporating images or videos that may exceed the input context windows of multimodal language models.
AI Agent with Tool Use. The second limitation of this work is that we do not consider tool use by LLMs when solving NPC problems. LLMs equipped with tool-using capabilities are typically referred to as AI agents (Wang et al., 2024a). Despite this limitation, the benchmark has significant potential to contribute to AI agent development by naturally encouraging tool use as problem difficulty scales. As problems become more complex, LLMs will increasingly need external tools to manage computational demands. This creates a natural progression toward agent capabilities, where models learn to decompose problems and leverage appropriate tools. Notably, we already observe code generation in models attempting difficult NPPC problems, which can be viewed as a form of tool creation, since these generated codes can be reused for future problem-solving.
Unstoppable RL vs. Ever-Scaling NP Problems. The rapid progress in LLM reasoning capabilities through reinforcement learning (RL) presents an interesting dynamic when considered alongside ever-scaling NPC problems. Recent work (Liu et al., 2025a; b) demonstrate that the reasoning data generated by NPC problems can improve the general reasoning capabilities of LLMs. As models like DeepSeek-R1 and OpenAI o1/o3-mini demonstrate significant reasoning improvements through RL techniques, NPPC provides a counterbalance by offering problems that can continuously scale in difficulty. This creates an adversarial paradigm: RL improves model reasoning and NPPC scales to maintain challenging (Zeng et al., 2025).
7 Conclusion